An article in The Wall Street journal implies that more than half of all Americans would prefer being given $100 rather than a day off from work. This statement is based on a survey conducted by American Express Incentive Services, in which 593 of 1,040 respondents indicated that they would rather have the $100 )Carlos Tejada, “ Work Week,” The Wall Street Journal, July 25, 2000, A1).
a. At the 0.05 level of significance, is there evidence based on the survey data that more than half of all Americans would rather have $100 than a day off of work?
b. Compute the p-value and interpret its meaning.

Answer by stanbon(75887)

An article in The Wall Street journal implies that more than half of all Americans would prefer being given $100 rather than a day off from work. This statement is based on a survey conducted by American Express Incentive Services, in which 593 of 1,040 respondents indicated that they would rather have the $100 )Carlos Tejada, “ Work Week,” The Wall Street Journal, July 25, 2000, A1).
a. At the 0.05 level of significance, is there evidence based on the survey data that more than half of all Americans would rather have $100 than a day off of work?
---
Ho: p <= 0.5
Ha: p > 0.5 (this is the claim)
----------------------------------
p-hat = 593/1040 = 0.5702
z(0.5702) = (0.5702-0.5)/[sqrt[0.5*0.5/1040] = 2.2639
p-value = P(z>2.2639) = 0.01179
------
Since the p-value is less than 5%, Fail to reject Ho.
Conclusion: The test results do not support the claim that p>0.5.
========================================================================
Cheers,
Stan H.
b. Compute the p-value and interpret its meaning.